OpenStudy (matlee):

Last year, Ann had $20,000 to invest. She invested some of it in an account that paid 5% simple interest per year, and she invested the rest in an account that paid 6% simple interest per year. After one year, she received a total of $1110 in interest. How much did she invest in each account?

OpenStudy (matlee):

I had to

TheSmartOne (thesmartone):

She has $20,000 in total to invest. She invests x dollars in an account that pays 5% simple interest per year She invests (20,000 - x) dollars in an account that pays 6% simple interest per year. After 1 year, she has made 1,100 dollars in total from the two investments.

TheSmartOne (thesmartone):

So, do you know the simple interest formula?

TheSmartOne (thesmartone):

@matlee

OpenStudy (matlee):

Hello! @TheSmartOne Nice to see you again, i had to walk my dog .

OpenStudy (matlee):

No

OpenStudy (matlee):

Nvm i guessed the numbers and got the right answer.

OpenStudy (matlee):

Please may someone help me with another questions instead

OpenStudy (matlee):

Keith works as a tutor for $8 an hour and as a waiter for $12 an hour. This month, he worked a combined total of 106 hours at his two jobs. Let t be the number of hours Keith worked as a tutor this month. Write an expression for the combined total dollar amount he earned this month.

TheSmartOne (thesmartone):

t = number of hours worked as a tutor 106 - t = number of hours worked as a waiter 8t + 12(106-t) is how much he will earn in total this month So just simplify it. First distribute the 12, and then add/subtract like terms. And that's your final answer :)

OpenStudy (matlee):

8t+1272-12t -4t+1272=0 +4t +4t 318.25 is that right?

OpenStudy (matlee):

It wants total earned in dollars

TheSmartOne (thesmartone):

correct :)

OpenStudy (matlee):

Thank you

OpenStudy (matlee):

Good rating!

OpenStudy (matlee):

May you help me with 1 more?

OpenStudy (matlee):

This is my last one

TheSmartOne (thesmartone):

Sure! Better to post it than ask because then even someone else can help instead of getting a confirmation from me :p

OpenStudy (anonymous):

Hello

OpenStudy (matlee):

No i had to go lol but thanks